A girl jogs around a horizontally circle with a constant speed. She travels one fourth of a revolution, a distance of 20.1 m along the circumference of the circle, in 4.7 s. The magnitude of her acceleration is ____ m/s2. Round your answer to the nearest tenth. Do not use scientific notation.
Answer #1

formula for centripetal accleration is

a= v^2/r

since v= 2pi r/T

= 2pi ( C/ 2pi)/T

= C/T

a= ( C/T)^2/r

--------------------------------------------------------------------------------------------------

( 2pi r/4) = 20.1 m

r= 12.8 m

a= ( 20.1/4.7)^2/12.8 m

=1.42 m/s^2